OpenStudy (anonymous):

What is the exact value of 4q^6 divided by the sqrt of q^21

OpenStudy (anonymous):

You want to get rid of the sort on the bottom so you multiply by sqrt q^21/sqrt q^21 ..which would be... [4q^6 / sqrt q^21] [(sqrt q^21)/(sqrt q^21)] = 4q^6(sqrt q^21) / q^21 = 4(sqrt q^21) / q^15

OpenStudy (anonymous):

can u draw the problem?

OpenStudy (anonymous):

\[\frac{ 4q^6 }{ \sqrt{q^{21}}}\]